a”Teddy cuts grass in the summer for $20 a lawn. His neighbor, Gavin is
very competitive and decides to cut grass as well but only charges $18 a
lawn. Their total combined revenue one summer was $348. Teddy made
$132 more than Gavin. How many lawns did they each cut?”

Answer by ikleyn(52788)

.
T + G = 348

T - G = 132

-------------------- Add the equations.


2T = 348 + 132 = 480  ======>  T = 480/2 = 240.  It is what Teddy earned.


G = 240 - 132 = 108.   It is what Gavin earned.


Teddy  cut  240/20 = 12 lawns.

Gavin  cut  108/18 =  6 lawns.


CHECK.   12*20 - 6*18 = 132.    ! Correct !

Solved, checked, explained and completed.